OpenStudy (anonymous):

Solve 1.005^n > 100, giving your answer correct to 3 sig figs. (Steps would be appreciated ^^)

OpenStudy (anonymous):

\[1.005^n>100\] \[n\log(1.005)>2\] \[n>\frac{2}{\log(1.005)}\]

OpenStudy (anonymous):

here i used log base ten because log base ten of 100 is 2. but you can use natural log if you like and get \[n>\frac{\ln(100)}{\ln(1.005)}\] using the same steps

OpenStudy (anonymous):

just to clarify when you are solving for a variable in the exponent, the first step is usually to take the log. that puts it on the ground floor. if you need even more steps worked out let me know

OpenStudy (anonymous):

That's cool. Thank you very much :)

OpenStudy (anonymous):

yw